OpenStudy (anonymous):

Please help! Conic Solutions? Name and Calculate the solutions for each system. 16x^2 - 3y^2 = -11 8x - y = -11 Thank you so much!

OpenStudy (anonymous):

Please help! Thank you!

OpenStudy (foreverandalways13):

@jdoe0001

OpenStudy (foreverandalways13):

welcome to openstudy!!

OpenStudy (anonymous):

The answer is supposed to be: (-1,3) and (-2,-5) but i dont know how Thanks!

OpenStudy (foreverandalways13):

np! im leah btw!

OpenStudy (anonymous):

Hi leah! im srini!

OpenStudy (foreverandalways13):

I fanned u so u can message me now...os works like facebook a little:)

OpenStudy (anonymous):

Oh okay ! awesome!

OpenStudy (foreverandalways13):

im here to help if u ever need anything!!!

OpenStudy (anonymous):

aww thank you! im still confused about this problem though like how they even got those answers :(

OpenStudy (foreverandalways13):

im not sure either :( oh well...

OpenStudy (foreverandalways13):

are u online school too?

OpenStudy (anonymous):

no.

OpenStudy (foreverandalways13):

oh...haha

OpenStudy (anonymous):

Are you supposed to be solving the equations?

OpenStudy (foreverandalways13):

no she figured it out...lol

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

I still didnt understand it.

OpenStudy (anonymous):

I think thats one part.

OpenStudy (anonymous):

but i dont understand the step by step things.

OpenStudy (jdoe0001):

have you done system of equations yet?

OpenStudy (anonymous):

yeah. but i keep getting weird answers

OpenStudy (jdoe0001):

ok... so, if you solve the bottom equation for "y" 8x - y = -11 y = 8x+11

OpenStudy (anonymous):

okay. then u plug it in right?

OpenStudy (jdoe0001):

so we use that \(\bf y = 8x+11\) in the above equation, yes

OpenStudy (anonymous):

but when i do that, i end up with a weird answer.

OpenStudy (jdoe0001):

well, let's see $$ 16x^2-3(8x+1)^2=-11\\ \text{using the binomial theorem}\\ (8x+1)^2 = 64x^2+16x+1 $$

OpenStudy (jdoe0001):

thus \(\bf 16x^2-3(64x^2+16x+1)=-11\\ 16x^2-192x^2-48x-3=-11\\ -176x^2-48x-3+11=0 \implies -176x^2-48x+8=0 \)

OpenStudy (anonymous):

okay but how does that equal those coordinates?

OpenStudy (anonymous):

thank you for helping me btw!

OpenStudy (jdoe0001):

from there we'd need to use the quadratic formula to get the "x" values keep in mind is a quadratic equation, degree is 2, so "x" will have 2 values

OpenStudy (jdoe0001):

\(\bf \text{quadratic formula}\\ x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\)

OpenStudy (jdoe0001):

wait.. lemme get common factors first

OpenStudy (jdoe0001):

\(\bf -176x^2-48x+8=0 \implies -8(22x^2+6x-1) = 0\\ \implies 22x^2+6x-1=0\)

OpenStudy (jdoe0001):

so we'll use the quadratic formula on that one

OpenStudy (anonymous):

idk why but im not getting the correct answer.

OpenStudy (jdoe0001):

$$\bf 22x^2+6x-1=0\\ x= \cfrac{ - 6 \pm \sqrt { 6^2 -4(22)(-1)}}{2(22)}\\ x= \cfrac{ - 6 \pm \sqrt { 36+88}}{44}\\ x= \cfrac{ - 6 \pm \sqrt { 124}}{44} \implies x= \cfrac{ - 6 \pm 2\sqrt {31}}{44}\\ $$

OpenStudy (jdoe0001):

$$\bf x= \cfrac{ - \cancel{6} \pm \cancel{2}\sqrt {31}}{\cancel{44}} \implies x= \cfrac{ - 3 \pm \sqrt {31}}{2} $$

OpenStudy (jdoe0001):

woops

OpenStudy (jdoe0001):

$$\bf x= \cfrac{ - \cancel{6} \pm \cancel{2}\sqrt {31}}{\cancel{44}} \implies x= \cfrac{ - 3 \pm \sqrt {31}}{22} $$

OpenStudy (jdoe0001):

and those are 2 values for "x"

OpenStudy (anonymous):

yeah but that doesnt match with the answers :(

OpenStudy (jdoe0001):

what answers do yo have as choices?

OpenStudy (jdoe0001):

I mean, that can be written in decimal form too

OpenStudy (anonymous):

he answer is supposed to be: (-1,3) and (-2,-5

OpenStudy (jdoe0001):

the one above is just the rational form

OpenStudy (jdoe0001):

lemme check

OpenStudy (anonymous):

how do i get it out of the rational form?

OpenStudy (jdoe0001):

to plot it?

OpenStudy (jdoe0001):

the graphs only meet at one point even though the parabola is symmetric I was thinking they'd meet at 2 points, the 2 "x" values, but they only meet at 1 near the parabola's vertex

OpenStudy (jdoe0001):

and is not the coordinates you have there

OpenStudy (jdoe0001):

to plot them, well just solve for "y" one equation will be y = 8x + 1 the other equation will be \(\bf y = \sqrt{\cfrac{16x^2+11}{3}} \)

OpenStudy (jdoe0001):

since they meet at one point, means one of the "x" values won't work for it

OpenStudy (jdoe0001):

anyhow, I just checked the decimal values, the positive root from the quadratic formula result is the good one

OpenStudy (jdoe0001):

which is in decimal x = 0.12

OpenStudy (jdoe0001):

and if you want the "y" just use that in the y = 8x + 1 => y = 8(0.12) +1

OpenStudy (jdoe0001):

the 2nd root of the quadratic equation is in decimal form -0.39 which yields the "y" coordinate of -2.12 so (-0.39, -2.12) BUT if you get the vertex of the parabola equation, you'll notice that point is below the vertex, so therefore it gets ruled out, and the positive root for 0.12 for "x" is the good one